OpenStudy (anonymous):

You have a 45, 45, 90 degrees triangle. The hypotenuse is 19, what are the lengths of sides A and B?

OpenStudy (anonymous):

they are the hypotenuse divided by the square root of 2

OpenStudy (anonymous):

since both angles are equal, both legs are equal if you call the lengths of the legs \(a\) and the hypotenuse \(h\) then by pythagoras \[a^2+a^2=h^2\] \[2a^2=h^2\] \[h=a\sqrt{2}\] or if you like \[a=\frac{h}{\sqrt{2}}\]

OpenStudy (anonymous):

thanks